# On endomorphism algebras of silting complexes over hereditary abelian categories

### v2

@@ 4. Idempotent quotients for several classes of algebras
    $\operatorname{\mathsf{mod}}\nolimits B$ can be viewed as a subcategory of $\operatorname{\mathsf{mod}}\nolimits A$. In particular, a $B$-module is indecomposable if and only if it is an indecomposable $A$-module.
@@
    $\operatorname{\mathsf{mod}}\nolimits B$ is closed under factor modules, submodules, kernels and cokernels in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    Given a morphism $f:X\to Y$ in $\operatorname{\mathsf{mod}}\nolimits B$, then $f$ is monomorphic (resp. epimorphic) in $\operatorname{\mathsf{mod}}\nolimits B$ if and only if it is monomorphic (resp. epimorphic) in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    A sequence $0\to X\to Y\to Z\to 0$ in $\operatorname{\mathsf{mod}}\nolimits B$ is exact if and only if it is exact when viewed as a sequence in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$, if $\operatorname{\mathsf{pd}}_{A}M\leq 1$, then $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\leq 1$.
@@
Assume $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, and we prove $\operatorname{\mathsf{pd}}_{A}M\geq 2$. Note that for an $A$-module in $(eA)^{\perp}=\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, it is $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$ if and only if it is a projective $A/\langle e\rangle$-module. Let $0\to L\to N\to M\to 0$ be a short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, where $N$ is a projective $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$-module. By Lemma 4.6, $N\cong P/PeA$ for some projective $A$-module $P$. By Lemma 4.8 (4), the above short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$ is also a short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A$. Considering the pull-back of morphisms $L\to N$ and $P\to N$, we have the following commutative diagram in $\operatorname{\mathsf{mod}}\nolimits A$:
@@
Since $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, $L$ is not $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Note that $PeA\in\operatorname{\mathsf{Fac}}eA$ and $L\in(eA)^{\perp}$. The first column in the above diagram is the canonical sequence of X. Hence, by Lemma 4.4, $L\cong X/XeA$. By Lemma 4.6 again, $X$ is non-projective in $\operatorname{\mathsf{mod}}\nolimits A$ since $L$ is not $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Hence $\operatorname{\mathsf{pd}}_{A}M\geq 2$, we obtain a contradiction. ∎
@@
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$, if $\operatorname{\mathsf{id}}_{A}M\leq 1$, then $\operatorname{\mathsf{id}}_{A/\langle e\rangle}M\leq 1$.
@@
For an indecomposable $A$-module $M$ in $(eA)^{\perp}$,
@@
We only prove (1); the proof of (2) is similar. Assume there is a path from an indecomposable module $X$ to $M$ in $\operatorname{\mathsf{ind}}A/\langle e\rangle$. Then this is also a path from $X$ to $M$ in $\operatorname{\mathsf{ind}}A$. Since $M\in\mathcal{L}_{A}$, then $\operatorname{\mathsf{pd}}_{A}X\leq 1$. By Proposition 4.9, we have $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}X\leq 1$. This means $M\in\mathcal{L}_{A/\langle e\rangle}$. ∎
@@
Remark 4.12.
The statement in Theorem 4.1 does not hold for tilted algebras and quasi-tilted algebras, while, in contrast, the result for special idempotents in [Z] remains valid for these two classes. Indeed, there is an algebra in the example from [BMR, Section 2.5] which is a tilted algebra, becoming strictly shod instead after taking a quotient by an idempotent.
@@ 5. Examples
We consider a class of algebras. Let $n\geq 2$, $2\leq k\leq n$. Denote by $A(n,k)$ the bound quiver algebra given by:
@@
Fix $n\geq 3$ and $2\leq k<n$. For each vertex $i$ of the quiver, we denote the corresponding simple module by $S(i)$, its projective cover by $P(i)$, and its injective envelop by $I(i)$. The Auslander-Reiten quiver of $A(n,k)$ has the following shape
@@
$S(1)=P(1)$ is projective. So $\operatorname{\mathsf{pd}}_{A(n,k)}S(1)=0$. For $1<i\leq k$, we have a short exact sequence $0\to S(i-1)\to P(i)\to S(i)\to 0$. By induction, $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=i-1$ for $1\leq i\leq k$. For $i>k$, we have a short exact sequence $0\to P(i-1)\to P(i)\to S(i)\to 0$. So $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=1$ for $k<i\leq n$. Hence $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. (By the same argument, one can check $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$.)
@@
Next we consider the endomorphism algebra $B=\operatorname{\mathsf{End}}_{A(n,k)}T$. Write $T=T_{1}\oplus\cdots\oplus T_{n}$ where $T_{i}=P(i)$ for $1\leq i\leq k-1$, $T_{k}=P(n)$, $T_{i}=I(i)$ for $k+1\leq i\leq n$. Observe the above Auslander-Reiten quiver and the positions of these $T_{i}$. We have the following facts:
  • For $1\leq i<j\leq n$, there is no path from $T_{j}$ to $T_{i}$ and hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{j},T_{i})=0$.
  • For $1\leq i\leq n-1$, there is only 1 path from $T_{i}$ to $T_{i+1}$ which is sectional. Composing the irreducible morphisms in this path, we obtain a nonzero morphism $f_{i}:T_{i}\to T_{i+1}$. Hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{i+1})$ is a 1-dimensional $k$-vector space spanned by $f_{i}$.
@@
  • For $k\leq i\leq n-2$, $i+2\leq j\leq n$, notice that $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{j})$ is a 1-dimensional $k$-vector space spanned by $f_{ij}=f_{j-1}f_{j-2}\cdots f_{i}$.
Denote by $e_{i}$ the primitive idempotent of $B$ corresponding to $T_{i}$. With the above observations, we have
@@
$e_{i+1}(\operatorname{\mathsf{rad}}B)e_{i}/e_{i+1}(\operatorname{\mathsf{rad}}B)^{2}e_{i}\cong kf_{i},~\text{ for }1\leq i\leq n-1,$
@@
In addition, $f_{2}f_{1}=0$, $f_{3}f_{2}=0$, $\cdots$, $f_{k}f_{k-1}=0$.
Hence $B$ is given by
@@
For $n\geq 2$ and $2\leq k\leq n$, we have $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. If $k<n$, there is a tilting $A(n,k)$-module $T$ such that $\operatorname{\mathsf{End}}_{A(n,k)}T\cong A(n,k+1)$. In particular, iteratively taking the endomorphism algebras of tilting modules, starting from a hereditary algebra of type $\mathbb{A}_{n}$, yields an algebra of global dimension n1n-1 after n2n-2 steps.
@@
It suffices to note that $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$ and $A(n,2)$ is a hereditary algebra of type $\mathbb{A}_{n}$. ∎
Let $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}X_{2}\to\cdots\xrightarrow{f_{s}}X_{s}$ be a path of irreducible morphisms in $\operatorname{\mathsf{ind}}A$. If $\tau X_{i+1}=X_{i-1}$ for some $1\leq i\leq s-1$, then $X_{i}$ is called a hook in this path (cf. [CL]).
We also have the following.
@@
with $\beta\alpha=0$. By Proposition 5.2 (2), $A\cong A(4,3)$ is quasi-tilted. Its Auslander-Reiten quiver is in the following shape
@@
We take $T=P(1)\oplus P(2)\oplus P(4)\oplus I(4)$, which is the Bongartz completion of $I(4)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}T$ is given by:
@@
We take $T=P(1)\oplus P(2)\oplus P(3)\oplus P(5)\oplus I(5)$, which is the Bongartz completion of $I(5)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}T$ is given by:
@@
with $\beta\alpha=0$, $\gamma\beta=0$, $\delta\gamma=0$. It is easy to check $\operatorname{\mathsf{pd}}_{B}S(3)=\operatorname{\mathsf{id}}_{B}S(3)=2$, and hence $B=\operatorname{\mathsf{End}}_{A}T$ is not shod.

### v3

@@ 4. Idempotent quotients for several classes of algebras
    The category $\operatorname{\mathsf{mod}}\nolimits B$ is identified with a full subcategory of $\operatorname{\mathsf{mod}}\nolimits A$ consisting of modules annihilated by $I$. In particular, a $B$-module is indecomposable if and only if it is indecomposable as an $A$-module.
@@
    $\operatorname{\mathsf{mod}}\nolimits B$ is closed under taking factor modules, submodules, kernels and cokernels in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    A morphism $f:X\to Y$ in $\operatorname{\mathsf{mod}}\nolimits B$ is a monomorphism (resp. epimorphism) if and only if it is a monomorphism (resp. epimorphism) when regarded as a morphism in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
    A sequence $0\to X\to Y\to Z\to 0$ of $B$-modules is exact in $\operatorname{\mathsf{mod}}\nolimits B$ if and only if it is exact in $\operatorname{\mathsf{mod}}\nolimits A$.
@@
Let $M$ be an indecomposable $A$-module belonging to $(eA)^{\perp}$. If $\operatorname{\mathsf{pd}}_{A}M\leq 1$, then $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\leq 1$.
@@
Conversely, assume $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, and we prove $\operatorname{\mathsf{pd}}_{A}M\geq 2$. Recall that for any $A$-module in the perpendicular category $(eA)^{\perp}\cong\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, being $\operatorname{\mathsf{Ext}}$-projective is equivalent to being a projective $A/\langle e\rangle$-module. Let $0\to L\to N\to M\to 0$ be a short exact sequence in $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$, where $N$ is a projective $\operatorname{\mathsf{mod}}\nolimits A/\langle e\rangle$-module. By Lemma 4.7, $N\cong P/PeA$ for some projective $A$-module $P$. According to Lemma 4.9 (4), this sequence remains exact when regarded as a sequence in $\operatorname{\mathsf{mod}}\nolimits A$. By taking the pull-back of the morphisms $L\to N$ and $P\to N$, we construct the following commutative diagram with exact rows and columns in $\operatorname{\mathsf{mod}}\nolimits A$:
@@
Since we assumed $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}M\geq 2$, the kernel $L$ can not be $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Observing the first column of the pull-back diagram, where $PeA\in\operatorname{\mathsf{Fac}}eA$ and $L\in(eA)^{\perp}$, we identify this as the canonical exact sequence for $X$. Hence, by Lemma 4.5, $L\cong X/XeA$. By Lemma 4.7 again, $X$ is non-projective in $\operatorname{\mathsf{mod}}\nolimits A$ since $L$ is not $\operatorname{\mathsf{Ext}}$-projective in $(eA)^{\perp}$. Hence $\operatorname{\mathsf{pd}}_{A}M\geq 2$. This provides the required contradiction. ∎
@@
Let $M$ be an indecomposable $A$-module belonging to $(eA)^{\perp}$. If $\operatorname{\mathsf{id}}_{A}M\leq 1$, then $\operatorname{\mathsf{id}}_{A/\langle e\rangle}M\leq 1$.
@@
Let $M$ be an indecomposable $A$-module belonging to $(eA)^{\perp}$.
@@
We restrict our attention to the proof of (1), as the argument for (2) follows by dual considerations. Assume there is a path from an indecomposable module $X$ to $M$ in $\operatorname{\mathsf{ind}}A/\langle e\rangle$. Then this is also a path from $X$ to $M$ in $\operatorname{\mathsf{ind}}A$. Since $M\in\mathcal{L}_{A}$, then $\operatorname{\mathsf{pd}}_{A}X\leq 1$. By Proposition 4.10, we have $\operatorname{\mathsf{pd}}_{A/\langle e\rangle}X\leq 1$. It follows that $M\in\mathcal{L}_{A/\langle e\rangle}$. ∎
@@ 5. Examples
We introduce a family of algebras $A(n,k)$ defined for $n\geq 2$ and $2\leq k\leq n$. Specifically, $A(n,k)$ is the bound quiver algebra $kQ/I$ given by the following quiver:
@@
Fix $n\geq 3$ and $2\leq k<n$. For each vertex $i$ of the quiver, we denote the corresponding simple module by $S(i)$, its projective cover by $P(i)$, and its injective envelope by $I(i)$. The Auslander-Reiten quiver of $A(n,k)$ has the following shape
@@
We determine the global dimension of $A(n,k)$ by computing the projective dimensions of its simple modules $S(i)$. First, consider the case for $1\leq i\leq k$. The simple module $S(1)$ coincides with the projective cover $P(1)$, so $\operatorname{\mathsf{pd}}_{A(n,k)}S(1)=0$. For $1<i\leq k$, the quadratic relations $\alpha_{j+1}\alpha_{j}=0$ induce the following short exact sequences:
$0\to S(i-1)\to P(i)\to S(i)\to 0.$
By induction on $i$, it follows that $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=i-1$ for $1\leq i\leq k$.
For $i>k$, we have a short exact sequence
$0\to P(i-1)\to P(i)\to S(i)\to 0.$
It follows that $\operatorname{\mathsf{pd}}_{A(n,k)}S(i)=1$ for all $k<i\leq n$. Hence $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. (By the same argument, one can check $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$.)
@@
We now investigate the structure of the endomorphism algebra $B=\operatorname{\mathsf{End}}_{A(n,k)}(T)$. Write $T=T_{1}\oplus\cdots\oplus T_{n}$, where $T_{i}=P(i)$ for $1\leq i\leq k-1$, $T_{k}=P(n)$, and $T_{i}=I(i)$ for $k+1\leq i\leq n$. By analyzing the distribution of these summands within the Auslander-Reiten quiver, we have the following facts:
  • For $1\leq i<j\leq n$, there are no paths from $T_{j}$ to $T_{i}$ and hence $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{j},T_{i})=0$.
  • For $1\leq i\leq n-1$, there exists a unique path in the AR quiver from $T_{i}$ to $T_{i+1}$ which is sectional. The composition of irreducible morphisms along this path yields a non-zero morphism $f_{i}:T_{i}\to T_{i+1}$. Accordingly, $\dim_{k}\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{i+1})=1$, with the space spanned by $f_{i}$.
@@
  • For $k\leq i\leq n-2$, $i+2\leq j\leq n$, the space $\operatorname{\mathsf{Hom}}_{A(n,k)}(T_{i},T_{j})$ is a 1-dimensional $k$-vector space spanned by $f_{ij}=f_{j-1}f_{j-2}\cdots f_{i}$.
Denote by $e_{i}$ the primitive idempotent of $B$ corresponding to $T_{i}$. Based on the homological observations above, the structure of the Jacobson radical $\operatorname{\mathsf{rad}}B$ is characterized as follows:
@@
$e_{i+1}(\operatorname{\mathsf{rad}}B)e_{i}/e_{i+1}(\operatorname{\mathsf{rad}}B)^{2}e_{i}\cong\operatorname{\mathsf{span}}\{f_{i}\},~\text{ for }1\leq i\leq n-1,$
@@
In addition, $f_{2}f_{1}=0$, $f_{3}f_{2}=0$, $\cdots$, $f_{k}f_{k-1}=0$. Consequently, the endomorphism algebra $B=\operatorname{End}_{A(n,k)}(T)$ is defined by the following bound quiver:
@@
For $n\geq 2$ and $2\leq k\leq n$, we have $\operatorname{\mathsf{gl.dim}}A(n,k)=k-1$. If $k<n$, there is a tilting $A(n,k)$-module $T$ such that $\operatorname{\mathsf{End}}_{A(n,k)}(T)\cong A(n,k+1)$. In particular, iteratively taking the endomorphism algebras of tilting modules, starting from a hereditary algebra of type $\mathbb{A}_{n}$, yields an algebra of global dimension n1n-1 after n2n-2 steps. Moreover, $A(n,k)$ is isomorphic to the endomorphism algebra of a tilting complex over a hereditary algebra of type $\mathbb{A}_{n}$.
@@
It suffices to note that $\operatorname{\mathsf{gl.dim}}A(n,n)=n-1$ and $A(n,2)$ is a hereditary algebra of type $\mathbb{A}_{n}$. For the last assertion, it suffices to observe that $A(n,k)$ is derived equivalent to $A(n,2)$, and the preimage of $T$ in ${\mathcal{D}}^{b}(A(n,2))$ is a tilting complex. ∎
Let $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}X_{2}\to\cdots\xrightarrow{f_{s}}X_{s}$ be a path of irreducible morphisms in $\operatorname{\mathsf{ind}}A$. If $\tau X_{i+1}=X_{i-1}$ for some $1\leq i\leq s-1$, then $X_{i}$ is called a hook in this path (see [CL]).
@@
with $\beta\alpha=0$. By Proposition 5.2 (2), $A\cong A(4,3)$ is quasi-tilted. Its Auslander-Reiten quiver exhibits the following structure:
@@
We take $T=P(1)\oplus P(2)\oplus P(4)\oplus I(4)$, which is the Bongartz completion of $I(4)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}(T)$ is given by:
@@
We take $T=P(1)\oplus P(2)\oplus P(3)\oplus P(5)\oplus I(5)$, which is the Bongartz completion of $I(5)$. The endomorphism algebra $B=\operatorname{\mathsf{End}}_{A}(T)$ is given by:
@@
with $\beta\alpha=0$, $\gamma\beta=0$, $\delta\gamma=0$. It is easy to check $\operatorname{\mathsf{pd}}_{B}S(3)=\operatorname{\mathsf{id}}_{B}S(3)=2$, and hence $B=\operatorname{\mathsf{End}}_{A}(T)$ is not shod.